OpenStudy (darkigloo):

integration help...given f(x)= x+1 for x<0 and cos(pix) for x=>0 find...

OpenStudy (darkigloo):

\[\int\limits_{-1}^{1} f(x) dx\]

zepdrix (zepdrix):

\[\Large\sf \int\limits\limits_{-1}^{1} f(x) dx\quad=\quad \int\limits_{-1}^{0} (x+1)\;dx+\int\limits_{0}^{1} \cos (\pi x)\;dx\]Understand what I did there? :o

OpenStudy (darkigloo):

ohh yes

zepdrix (zepdrix):

Understand how to finish it? ^^

OpenStudy (darkigloo):

yup i got it. thank you! :)